### v1

@@ 6 Numerical experiments
Experiment 6.1.
We compute the $10$ singular triplets of $A=\mathrm{Alemdar}$ with the singular values closest to $\tau=67.2$, which are the largest ones and clustered.
Refer to captionRefer to caption
Figure 1: Computing $10$ singular triplets of $A=\mathrm{Alemdar}$ with $\tau=67.2$.
Both IPJDSVD and JDSVD succeeded in computing the $10$ desired singular triplets of $A$ in the correct order. Figure 1 (left) draws the outer iterations used by IPJDSVD and JDSVD to compute each singular triplet, and Figure 1 (right) depicts the average inner iterations per outer iteration of IPJDSVD and JDSVD to compute each desired singular triplet. It is seen from the figure that IPJDSVD used the same outer iterations for the first, second, fourth, sixth and eighth singular triplets as JDSVD, but consumed fewer outer iterations than JDSVD for the fifth, seventh, ninth and last ones, showing the effectiveness of the new thick-restart. Regarding the overall efficiency, IPJDSVD, on average, saved $104$ inner iterations for MINRES to solve the preconditioned correction equation at each outer iteration. IPJDSVD totally consumed $54$ outer iterations, $7433$ inner iterations and $1.83$ seconds, while JDSVD used $59$ outer iterations, $12472$ inner iterations and $2.53$ seconds. IPJDSVD outperforms JDSVD considerably, and reduced $8.47\%$ of the outer iterations, $40.40\%$ of the inner iterations, and $27.76\%$ of the CPU time.
Experiment 6.2.
[…]
Refer to captionRefer to caption
[…]
From Figure 2, we observe that for the first six singular triplets of $A$, IPJDSVD used almost the same outer iterations as JDSVD did; but for the last four desired singular triplets of $A$, IPJDSVD was much faster and used substantially fewer outer iterations, illustrating that the new thick-restart can play a crucial role. More importantly, for each desired singular triplet, the preconditioned MINRES was much faster than the unpreconditioned one.
Precisely, IPJDSVD consumed $73$ outer iterations, $7051$ inner iterations and $3.74$ seconds, while JDSVD used $87$ outer iterations, $12225$ inner iterations and $5.72$ seconds. The former outmatched the latter considerably by reducing $16.09\%$ of the outer iterations, $42.32\%$ of the inner iterations, and $34.61\%$ of the CPU time.
Experiment 6.3.
We compute $10$ singular triplets of four test matrices in Table 1 with given targets $\tau$’s, where the desired singular values are not very interior.
Table 2: Results on Experment 6.3.
@@
$I_{\mathrm{out}}$ $I_{\mathrm{in}}$ $T_{\rm cpu}$ $I_{\mathrm{out}}$ $I_{\mathrm{in}}$ $T_{\rm cpu}$
cat_ears_3_4 4.9 74 6627 2.19 84 11219 2.76
epb2 3.3 53 5176 3.68 61 9688 5.09
relat8 16 64 7510 83.2 70 12533 110
Rucci1 6.7 55 5302 331 63 8039 415
Table 2 lists the results. IPJDSVD and JDSVD succeeded in computing the desired singular triplets of all the test matrices. Regarding the outer convergence, we observe from Table 2 that, for all the test matrices, IPJDSVD converged faster than JDSVD, confirming the higher effectiveness of the new thick-restart. Meanwhile, for all the test matrices, IPJDSVD reduced more than $34\%$ of the total inner iterations and more than $20\%$ of the CPU time by JDSVD.
Obviously, for these matrices, IPJDSVD outperforms JDSVD considerably.
Experiment 6.4.
We compute $\ell=10$ very interior and clustered singular triplets of the nine matrices in Table 1 with given $\tau$’s.
Table 3: Results on Experiment 6.4.
$A$ $\tau$ Algorithm: IPJDSVD Algorithm: JDSVD
$I_{\mathrm{out}}$ $I_{\mathrm{in}}$ $T_{\rm cpu}$ $I_{\mathrm{out}}$ $I_{\mathrm{in}}$ $T_{\rm cpu}$
dw8192 109.7 72 4845 2.06 171 11469 3.02
hep-th 11 63 20253 7.34 60 31334 10.3
ca-HepTh 8 65 22019 8.65 61 37117 12.8
California 6 67 30252 9.00 67 46918 12.2
inlet 5.6 75 23565 10.3 85 42909 16.2
garon2 23 79 47190 21.5 91 81969 34.0
e40r0100 13 72 82225 51.1 66 116456 60.8
tomographic1 5.6 57 11227 31.4 63 19629 47.7
degme 28.6 69 18102 565 61 29744 689
[…]
@@ 7 Conclusions
We have analyzed the convergence of MINRES for the correction equations in JDSVD. Based on the results obtained, we have proposed an inner preconditioning procedure for the correction equation and derived the preconditioned correction equation, whose solution is proved to have the same accuracy as that to the standard correction equation. We have extended the convergence results on MINRES for the standard correction equation to the preconditoned MINRES, and shown that MINRES for solving the preconditioned correction equation can converge much faster than it does for solving the standard correction equation.
We have presented a practical strategy that dynamically selects approximate singular triplets from current available ones and forms the preconditioned correction equation. Finally, we have developed a new thick-restart IPJDSVD algorithm with deflation and purgation to compute several singular triplets. Numerical experiments have illustrated that the new thick-restart IPJDSVD algorithm outperforms the standard thick-restart JDSVD algorithm considerably, in terms of inner and outer iterations as well as CPU time.
@@
The two authors declare that they have no financial interests, and they read and approved the final manuscript.

### v2

@@ 6.1 A comparison of IPJDSVD and JDSVD
Table 2: Results on computing the $10$ singular triplets with given targets.
@@
$I_{\mathrm{out}}$ MVs $T_{\rm cpu}$ $I_{\mathrm{out}}$ MVs $T_{\rm cpu}$
Alemdar 67.2 53 15998 1.63 59 24928 2.22
cat_ears_3_4 4.9 72 15262 2.32 86 22974 2.70
epb2 3.3 51 10968 3.50 60 19396 4.68
relat8 16 64 17036 77.2 70 25276 89.6
bcsstk33 56 75 15176 4.47 84 23680 6.12
dw8192 109.7 68 9952 1.59 227 26404 3.06
garon2 23 81 99810 22.2 91 161248 34.2
tomographic1 5.6 55 23494 28.7 61 36832 36.3
We computed $10$ singular triplets of eight test matrices in Table 1 with given targets $\tau$, where the desired singular values of the first four matrices are not very interior, and those of the last four matrices are interior and clustered ones. Both algorithms succeeded in computing all the desired singular triplets of all the test matrices and Table 2 lists the results. Regarding $I_{\operatorname{out}}$, we observe that IPJDSVD converged faster for all the test matrices and particularly much faster for “cat_ears_3_4” and “dw8192” and “garon2” than JDSVD, confirming the higher effectiveness of the new thick-restart. Regarding the overall efficiency, Fig. 1 depicts the average MVs per outer iteration of IPJDSVD and JDSVD to compute each desired singular triplet of “Alemdar” and “bcsstk33”. We can see that IPJDSVD, on average, saved $121$ and $80$ MVs at each outer iteration from JDSVD for the former and latter matrices, respectively, huge improvements. For all the test matrices, we can see from Table 2 that IPJDSVD reduced more than $32.74\%$ of MVs and more than $13.79\%$ of the CPU time by JDSVD.
Obviously, for these matrices, IPJDSVD is substantially superior to JDSVD.
6.2 A comparison of IPJDSVD_HYBRID and PRIMME_SVDS
We compare IPJDSVD_HYBRID with PRIMME_SVDS. For PRIMME_SVDS, we took the method in the first stage as the default ‘DYNAMIC’, dynamically switching between JDQMR and the GD+k method to minimize the time, and in the second stage we took the recommended ‘JDQMR_ETol’. For both algorithms, we took the maximum and minimum subspace dimensions $k_{\max}=30$ and $k_{\min}=3$ and the initial $u_{0}$ and $v_{0}$ as the same $m$- and $n$-dimensional random vectors generated by the Matlab function randn. All the other parameters in the two algorithms were set by default.
Table 3: Results on computing the $10$ largest singular triplets.
$A$ IPJDSVD_HYBRID PRIMME_SVDS IRLBD
$I_{\mathrm{out}}$ MVs $T_{\rm cpu}$ MVs $T_{\rm cpu}$ MVs $T_{\rm cpu}$
Alemdar 77 1804 2.05e-1 1858 1.27e-1 472 1.05e-1
3elt 103 2196 2.64e-1 2362 1.41e-1 774 1.58e-1
plddb 81 7680 3.46e-1 3806 1.12e-1 122820 20.1
bas1lp 46 522 3.56e-1 484 1.26e-1 86 6.44e-2
G66 83 2040 2.73e-1 2354 1.68e-1 726 1.72e-1
stat96v5 82 1376 2.82e-1 1306 2.10e-1 250 1.56e-1
atmosmodl 137 11334 6.05e+2 12926 2.02e+2 100184 3.97e+3
cage15 136 5364 1.02e+3 5528 4.99e+2 7168 1.12e+3
Table 4: Results on computing the $10$ smallest singular triplets.
$A$ IPJDSVD_HYBRID PRIMME_SVDS PRIMME_SVDS_J
$I_{\mathrm{out}}$ MVs $T_{\rm cpu}$ MVs $T_{\rm cpu}$ MVs $T_{\rm cpu}$
Alemdar 81 191462 10.6 312848 27.4 247150 69.0
3elt 190 117838 6.39 157086 12.2 146348 84.2
plddb 85 26484 6.59e-1 36328 1.77 37478 2.19
bas1lp 72 337718 69.2 416200 130 537622 189
G66 324 237528 29.5 81000005 9.19e+3 321304 77.9
stat96v5 171 16852 1.78 24972 5.78 24008 6.31
atmosmodl 62 625448 8.41e+3 382610 9.00e+3 313958 1.64e+4
cage15 90 4034 727 5316 722 2882 3.42e+3
We computed the $10$ largest and smallest singular triplets of the eight matrices in Table 1 with the outer stopping tolerance $\textit{tol}=10^{-12}$ by setting the target $\tau=\|A\|_{e}$ and 00 for IPJDSVD_HYBRID and the Sigma as ‘L’ and ‘S’ for PRIMME_SVDS, respectively. Table 4 reports the results. For the largest singular triplets, we also report the results obtained by the implicitly restarted LBD algorithm (IRLBD) [14] with the same parameters. For the smallest singular triplets, we also report the results computed by PRIMME_SVDS with the block Jacobi preconditioners (PRIMME_SVDS_J) with block size $10$ and $100$ for the first six and last two matrices, respectively.
When computing the largest singular triplets, IPJDSVD_HYBRID and PRIMME
_SVDS consumed comparable MVs and even runtime for all the matrices but “plddb”, and IRLBD outperformed them for five of the eight matrices by using significantly fewer MVs and less CPU time. For “plddb”, “atmosmodl” and “cage15”, the former two algorithms were substantially superior to IRLBD. We notice that IRLBD consumed much more MVs than PRIMME_SVDS for “plddb” and “atmosmodl”, and IPJDSVD_HYBRID did similarly for “plddb”. This is because some of the desired largest singular values are multiple, causing that IRLBD and IPJDSVD_HYBRID perform poorly while PRIMME_SVDS with block size bigger than or equal to the maximum multiplicity is able to compute multiple singular values.
As a whole, for computing the simple largest singular triplets, IPJDSVD_HYBRID is at least competitive with PRIMME_SVDS; IRLBD is generally preferred if the largest singular values are not clustered, otherwise the former algorithms may be advantageous.
For the computation of smallest singular triplets, except “atmosmodl” for which IPJDSVD_HYBRID used nearly as twice MVs as PRIMME_SVDS, for the other seven matrices, IPJDSVD_HYBRID performed considerably better than the latter regarding total MVs and CPU time; on average, IPJDSVD_HYBRID saves $25\%$ of MVs for each matrix, and it even used much less runtime than PRIMME_SVDS. Particularly, for “G66”, IPJDSVD_HYBRID converged fast and computed all the desired singular triplets correctly, while PRIMME_SVDS failed to converge when the maximum $N^{2}$ MVs were performed and only output five converged singular triplets.
When block Jacobi preconditioners were exploited in PRIMME_SVDS for the involved correction equations, we see they only accelerated the convergence of the inner iterations for “G66” and “cage15” dramatically and did not help for the other matrices in terms of CPU time and almost all the MVs. Overall, IPJDSVD_HYBRID outmatched PRIMME_SVDS_J considerably in terms of both MVs and CPU time, except for the last two matrices where IPJDSVD_HYBRID used more MVs. Therefore, IPJDSVD_HYBRID outperforms PRIMME_SVDS and PRIMME_SVDS_J greatly for computing the simple smallest singular triplets.
@@ 7 Conclusions
We have analyzed the convergence of MINRES for the correction equations in JDSVD, and proposed an inner preconditioning procedure for the correction equation and derived the preconditioned correction equation, whose solution has been proven to expand the subspaces as effectively as that to the standard correction equation. The resulting IPJDSVD method is essentially a new JDSVD method with a new correction equation solved at each outer iteration. We have established convergence results on the preconditioned MINRES, and shown that it can converge much faster than MINRES does for solving the standard correction equation.
We have presented a practical strategy that dynamically selects approximate singular triplets from current available ones and forms the preconditioned correction equation. Finally, we have developed a new thick-restart IPJDSVD algorithm with deflation and purgation to compute several singular triplets. Numerical experiments have illustrated that the new thick-restart IPJDSVD algorithm outperforms the standard thick-restart JDSVD algorithm considerably in terms of the outer iterations, MVs and CPU time; in the meantime, IPJDSVD_HYBRID is at least competitive with PRIMME_SVD in [27] for computing the largest singular triplets and outperforms the latter substantially for computing the smallest singular triplets in terms of MVs, provided that the desired singular values are simple.
@@
The two authors declare that they have no conflict of interest, and they read and approved the final manuscript.